A person deposits $1500 in a savings account that pays 3% annual interest compounded yearly. At the end of 12 years, what will be the balance in the savings account if no additional deposits or withdrawals are made on the account?

Answer by jim_thompson5910(35256)

We'll need the formula A = P*(1+r/n)^(n*t) where

A = final amount (ie future value)
P = initial amount which is the deposit
r = interest rate
n = compounding frequency
t = number of years

In this problem, we know

A = unknown (we're finding this value)
P = 1500
r = 0.03 (3% = 3/100 = 0.03)
n = 1 (compounding annually, so once per year)
t = 12 (the money sits in the account for 12 years)

So we take those values and substitute them into the formula. Afterwards, we use a calculator to get

A = P*(1+r/n)^(n*t)

A = 1500*(1+0.03/1)^(1*12)

A = 2,138.64133026927

A = 2,138.64

So $2,138.64 is the balance at the end of the 12 years.